OpenStudy (anonymous):

help me please Which system of equations does not have a real solution? y = x **2 + 3x - 5 and x + y = -10 y = x **2 + 3x - 5 and 4x + 5y = 20 y = x **2+ 3x- 5 and x + y = -9

OpenStudy (anonymous):

please be nice and someone help me

OpenStudy (anonymous):

which do you think

OpenStudy (anonymous):

@LmL

OpenStudy (anonymous):

I dont know :(

OpenStudy (anonymous):

damn my sleek-feathered one?<gone help you

OpenStudy (anonymous):

pretty asf

OpenStudy (anonymous):

told you bruh

OpenStudy (anonymous):

: x = 2y - 2 and 3x - y = 14 3(2y-2)-y=14 substitute x in first for x in second and distribute the 3 6y-6-y=14 combine like terms 5y=20 divide by 5 y=4 (only one with y=4 is the first, so that one. But I'll solve for x so you know how to) x=2(4)-2 x=8-2 x=6 (6,4) 2m - 3n = 6 and m = 3 - 2n 2(3-2n)-3n=6 substitute m in second for m in first and distribute the 2 6-4n-3n=6 combine like terms -7n=0 divide by -7 n=0 (again only one with n=0 is second, but I will solve for m) m=3-2(0) m=3 (3,0) 4x + y = 4 and y = 4 - 4x 4x+4-4x=4 substitute y in second for y in first 4=4 infinitely many solutions x = 4y and 2x + 3y = 44 2(4y)+3y=44 8y+3y=44 11y=44 y=4 x=4(4) x=16 (16,4) a + b = 4 and 2a + 3b = 11 b=-a+4 2a+3(-a+4)=11 2a-3a+12=11 -a+12=11 -a=-1 a=1 1+b=4 b=3 (1,3) y = 1 - 5x and 5x + y = 8 5x+1-5x=8 0=7 no solution 3a - b = 11 and 2a + 3b = 0 -b=-3a+11 b=3a-11 2a+3(3a-11)=0 2a+9a-33=0 11a=33 a=3 2(3)+3b=0 6+3b=0 3b=-6 b=-2 (3,-2)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Thanks :)